# : Generalizing Large Language Models for E-commerce

### v1

@@ Appendix F Complete Experimental Results
In this table, ``IND" indicates in-domain evaluation; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; ``NPR" is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | ||||||||||
| Acc | M-Rec | M-Pre | M-F1 | #failed | Acc | M-Rec | M-Pre | M-F1 | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.595 | 0.575 | 0.544 | 0.516 | 0 | 0.556 | 0.586 | 0.544 | 0.510 | 0 | |
| Gemini Pro | 0.609 | 0.521 | 0.453 | 0.470 | 2 | 0.572 | 0.511 | 0.444 | 0.454 | 1 | ||
| Claude 2.1 | 0.375 | 0.510 | 0.474 | 0.415 | 2 | 0.328 | 0.466 | 0.447 | 0.369 | 1 | ||
| Llama-2 13B-chat | 0.406 | 0.188 | 0.191 | 0.188 | 0 | 0.384 | 0.179 | 0.180 | 0.178 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.633 | 0.532 | 0.551 | 0.470 | 0 | 0.594 | 0.531 | 0.494 | 0.438 | 0 | ||
| E-commerce LLM | EcomGPT | 0.191 | 0.362 | 0.341 | 0.188 | 6 | 0.196 | 0.375 | 0.336 | 0.178 | 13 | |
| SoTA task-specific model | BERTweet | 0.733 | 0.503 | 0.530 | 0.511 | 0 | 0.729 | 0.507 | 0.524 | 0.513 | 0 | |
| DeBERTaV3 | 0.768 | 0.567 | 0.607 | 0.573 | 0 | 0.764 | 0.565 | 0.591 | 0.567 | 0 | ||
| P5 | 0.611 | 0.199 | 0.157 | 0.156 | 0 | 0.620 | 0.200 | 0.124 | 0.153 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.783 | 0.619 | 0.618 | 0.612 | 0 | 0.770 | 0.604 | 0.601 | 0.600 | 0 | |
| Llama-2 13B-chat | 0.791 | 0.616 | 0.641 | 0.616 | 0 | 0.781 | 0.627 | 0.645 | 0.629 | 0 | ||
| Llama-2 7B-chat | 0.790 | 0.620 | 0.652 | 0.634 | 0 | 0.769 | 0.583 | 0.599 | 0.589 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.801 | 0.643 | 0.676 | 0.655 | 0 | 0.789 | 0.619 | 0.650 | 0.632 | 0 | ||
| Flan-T5 XL | 0.771 | 0.645 | 0.638 | 0.620 | 0 | 0.743 | 0.594 | 0.592 | 0.582 | 0 | ||
| Phi-2 | 0.779 | 0.611 | 0.618 | 0.608 | 0 | 0.754 | 0.576 | 0.594 | 0.583 | 0 | ||
| Generalist | Flan-T5 XXL | 0.797 | 0.629 | 0.646 | 0.628 | 0 | 0.787 | 0.619 | 0.624 | 0.619 | 0 | |
| Llama-2 13B-chat | 0.796 | 0.641 | 0.661 | 0.648 | 0 | 0.785 | 0.621 | 0.638 | 0.629 | 0 | ||
| Llama-2 7B-chat | 0.768 | 0.579 | 0.589 | 0.580 | 0 | 0.776 | 0.599 | 0.626 | 0.606 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.781 | 0.630 | 0.654 | 0.639 | 0 | 0.784 | 0.630 | 0.653 | 0.640 | 0 | ||
| Flan-T5 XL | 0.782 | 0.654 | 0.655 | 0.648 | 0 | 0.753 | 0.604 | 0.598 | 0.598 | 0 | ||
| Phi-2 | 0.780 | 0.588 | 0.619 | 0.596 | 0 | 0.758 | 0.552 | 0.590 | 0.565 | 0 | ||
@@
In this table, ``IND" and ``OOD" indicates in-domain evaluation and out-of-domain evaluation, respectively; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric ``Acc'', ``M-Rec", ``M-Pre", and ``M-F1" represents accuracy, macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
@@
In this table, ``IND" indicates in-domain evaluation; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric ``M-Rec", ``M-Pre", and ``M-F1" represents macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
@@
In this table, ``IND" indicates in-domain evaluation; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; ``NPR" is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | ||||||||||||||
| Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.623 | 0.550 | 0.791 | 0.649 | 0.749 | 0.491 | 5 | 0.641 | 0.555 | 0.876 | 0.680 | 0.828 | 0.460 | 2 | |
| Gemini Pro | 0.542 | 0.371 | 0.797 | 0.506 | 0.837 | 0.435 | 13 | 0.543 | 0.410 | 0.844 | 0.552 | 0.834 | 0.393 | 6 | ||
| Claude 2.1 | 0.424 | 0.177 | 0.671 | 0.280 | 0.850 | 0.375 | 90 | 0.384 | 0.146 | 0.769 | 0.245 | 0.904 | 0.326 | 73 | ||
| Llama-2 13B-chat | 0.534 | 0.608 | 0.638 | 0.623 | 0.406 | 0.375 | 0 | 0.541 | 0.605 | 0.688 | 0.644 | 0.401 | 0.317 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.522 | 0.539 | 0.647 | 0.588 | 0.493 | 0.383 | 120 | 0.537 | 0.523 | 0.725 | 0.608 | 0.567 | 0.352 | 126 | ||
| E-commerce LLM | EcomGPT | 0.318 | 0.051 | 0.283 | 0.086 | 0.779 | 0.322 | 254 | 0.286 | 0.085 | 0.403 | 0.140 | 0.726 | 0.266 | 245 | |
| SoTA task-specific model | BERT | 0.749 | 0.970 | 0.726 | 0.830 | 0.368 | 0.877 | 0 | 0.803 | 0.831 | 0.876 | 0.853 | 0.742 | 0.668 | 0 | |
| DeBERTaV3 | 0.504 | 0.310 | 0.769 | 0.441 | 0.839 | 0.413 | 0 | 0.501 | 0.345 | 0.826 | 0.487 | 0.841 | 0.370 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.749 | 0.875 | 0.763 | 0.815 | 0.531 | 0.712 | 0 | 0.799 | 0.889 | 0.830 | 0.859 | 0.602 | 0.713 | 0 | |
| Llama-2 13B-chat | 0.801 | 0.919 | 0.797 | 0.854 | 0.597 | 0.811 | 0 | 0.838 | 0.924 | 0.852 | 0.887 | 0.650 | 0.797 | 0 | ||
| Llama-2 7B-chat | 0.741 | 0.889 | 0.749 | 0.813 | 0.485 | 0.718 | 0 | 0.761 | 0.864 | 0.802 | 0.832 | 0.535 | 0.644 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.821 | 0.896 | 0.834 | 0.864 | 0.692 | 0.794 | 0 | 0.835 | 0.891 | 0.872 | 0.881 | 0.713 | 0.749 | 0 | ||
| Flan-T5 XL | 0.693 | 0.684 | 0.802 | 0.738 | 0.708 | 0.565 | 0 | 0.707 | 0.682 | 0.862 | 0.762 | 0.761 | 0.523 | 0 | ||
| Phi-2 | 0.765 | 0.942 | 0.751 | 0.835 | 0.460 | 0.820 | 0 | 0.781 | 0.950 | 0.779 | 0.856 | 0.411 | 0.791 | 0 | ||
| Generalist | Flan-T5 XXL | 0.774 | 0.910 | 0.773 | 0.836 | 0.540 | 0.776 | 0 | 0.814 | 0.914 | 0.832 | 0.871 | 0.596 | 0.760 | 0 | |
| Llama-2 13B-chat | 0.808 | 0.864 | 0.838 | 0.851 | 0.711 | 0.752 | 0 | 0.818 | 0.862 | 0.872 | 0.867 | 0.723 | 0.705 | 0 | ||
| Llama-2 7B-chat | 0.817 | 0.921 | 0.814 | 0.864 | 0.638 | 0.824 | 0 | 0.836 | 0.931 | 0.845 | 0.886 | 0.627 | 0.807 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.797 | 0.880 | 0.814 | 0.846 | 0.654 | 0.759 | 0 | 0.832 | 0.885 | 0.872 | 0.878 | 0.717 | 0.740 | 0 | ||
| Flan-T5 XL | 0.765 | 0.888 | 0.774 | 0.827 | 0.553 | 0.741 | 0 | 0.819 | 0.891 | 0.852 | 0.871 | 0.662 | 0.735 | 0 | ||
| Phi-2 | 0.794 | 0.897 | 0.801 | 0.846 | 0.616 | 0.777 | 0 | 0.823 | 0.937 | 0.828 | 0.879 | 0.573 | 0.807 | 0 | ||
@@
In this table, ``IND" and ``OOD" indicates in-domain evaluation and out-of-domain evaluation, respectively; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; ``NPR" is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | ||||||||||
| R | P | F | BLEURT | #failed | R | P | F | BLEURT | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.847 | 0.869 | 0.858 | 0.280 | 0 | 0.852 | 0.868 | 0.860 | 0.283 | 0 | |
| Gemini Pro | 0.844 | 0.867 | 0.855 | 0.269 | 0 | 0.847 | 0.866 | 0.856 | 0.264 | 0 | ||
| Claude 2.1 | 0.848 | 0.835 | 0.841 | 0.314 | 0 | 0.851 | 0.833 | 0.842 | 0.325 | 0 | ||
| Llama-2 13B-chat | 0.845 | 0.780 | 0.811 | 0.261 | 0 | 0.845 | 0.775 | 0.808 | 0.260 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.850 | 0.856 | 0.853 | 0.288 | 0 | 0.852 | 0.851 | 0.851 | 0.290 | 0 | ||
| E-commerce LLM | EcomGPT | 0.675 | 0.665 | 0.669 | 0.290 | 0 | 0.729 | 0.716 | 0.722 | 0.296 | 0 | |
| SoTA task-specific model | GPT-4 Turbo | 0.847 | 0.869 | 0.858 | 0.280 | 0 | 0.852 | 0.868 | 0.860 | 0.283 | 0 | |
| Task-specific | Flan-T5 XXL | 0.822 | 0.864 | 0.842 | 0.310 | 0 | 0.824 | 0.865 | 0.843 | 0.302 | 0 | |
| Llama-2 13B-chat | 0.824 | 0.861 | 0.841 | 0.309 | 0 | 0.821 | 0.860 | 0.840 | 0.289 | 0 | ||
| Llama-2 7B-chat | 0.822 | 0.861 | 0.841 | 0.301 | 0 | 0.820 | 0.861 | 0.840 | 0.289 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.823 | 0.860 | 0.841 | 0.310 | 0 | 0.823 | 0.861 | 0.842 | 0.298 | 0 | ||
| Flan-T5 XL | 0.823 | 0.864 | 0.843 | 0.320 | 0 | 0.824 | 0.864 | 0.843 | 0.307 | 0 | ||
| Phi-2 | 0.817 | 0.855 | 0.835 | 0.283 | 0 | 0.817 | 0.856 | 0.835 | 0.270 | 0 | ||
| Generalist | Flan-T5 XXL | 0.824 | 0.865 | 0.844 | 0.224 | 0 | 0.823 | 0.864 | 0.843 | 0.206 | 0 | |
| Llama-2 13B-chat | 0.823 | 0.861 | 0.841 | 0.215 | 0 | 0.822 | 0.861 | 0.841 | 0.195 | 0 | ||
| Llama-2 7B-chat | 0.822 | 0.860 | 0.840 | 0.208 | 0 | 0.819 | 0.859 | 0.838 | 0.188 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.822 | 0.864 | 0.842 | 0.213 | 0 | 0.821 | 0.862 | 0.840 | 0.194 | 0 | ||
| Flan-T5 XL | 0.823 | 0.864 | 0.843 | 0.227 | 0 | 0.824 | 0.865 | 0.844 | 0.211 | 0 | ||
| Phi-2 | 0.823 | 0.861 | 0.842 | 0.222 | 0 | 0.821 | 0.859 | 0.840 | 0.198 | 0 | ||
@@
In this table, ``IND" and ``OOD" indicates in-domain evaluation and out-of-domain evaluation, respectively; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metrics ``R", ``P", ``F" and ``BLEURT" are detailed in TableΒ A1. On each task, the best performance is in bold. Note that we use GPT-4 Turbo as the SoTA task-specific model in this task.
| Model | ||||||||
| IND | OOD | IND | ||||||
| HR@1 | #failed | HR@1 | #failed | NDCG | #failed | |||
| General-purpose LLMs | GPT-4 Turbo | 0.387 | 0 | 0.198 | 0 | 0.875 | 14 | |
| Gemini Pro | 0.269 | 2 | 0.116 | 3 | 0.821 | 52 | ||
| Claude 2.1 | 0.066 | 34 | 0.036 | 42 | 0.821 | 26 | ||
| Llama-2 13B-chat | 0.056 | 0 | 0.050 | 0 | 0.815 | 0 | ||
| Mistral-7B-Instruct-v0.2 | 0.164 | 1 | 0.108 | 0 | 0.842 | 4 | ||
| E-commerce LLM | EcomGPT | 0.042 | 344 | 0.023 | 391 | 0.000 | 1000 | |
| SoTA task-specific model | gSASRec / BERT | 0.249 | 0 | 0.065 | 0 | 0.839 | 0 | |
| Recformer / DeBERTaV3 | 0.265 | 0 | 0.081 | 0 | 0.859 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.467 | 0 | 0.252 | 0 | 0.881 | 0 | |
| Llama-2 13B-chat | 0.518 | 0 | 0.263 | 0 | 0.879 | 0 | ||
| Llama-2 7B-chat | 0.517 | 0 | 0.228 | 0 | 0.867 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.535 | 0 | 0.268 | 0 | 0.883 | 0 | ||
| Flan-T5 XL | 0.436 | 0 | 0.226 | 0 | 0.875 | 0 | ||
| Phi-2 | 0.413 | 5 | 0.219 | 10 | 0.858 | 0 | ||
| Generalist | Flan-T5 XXL | 0.512 | 0 | 0.262 | 0 | 0.885 | 0 | |
| Llama-2 13B-chat | 0.526 | 0 | 0.273 | 0 | 0.870 | 0 | ||
| Llama-2 7B-chat | 0.517 | 0 | 0.261 | 0 | 0.868 | 0 | ||
| Mistral-7B Instruct-v0.2 | 0.542 | 0 | 0.280 | 0 | 0.876 | 0 | ||
| Flan-T5 XL | 0.463 | 0 | 0.256 | 0 | 0.868 | 0 | ||
| Phi-2 | 0.479 | 5 | 0.241 | 8 | 0.870 | 0 | ||
@@
In this table, ``IND" and ``OOD" indicates in-domain evaluation and out-of-domain evaluation, respectively; ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metrics ``HR@1" and ``NDCG" are detailed in AppendixΒ A. On each task, the best performance is in bold.
Appendix G OOD Evaluation

### v2

@@ Appendix G Complete Experimental Results
In this table, βINDβ indicates in-domain evaluation; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; βNPRβ is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||
| Acc | M-Rec | M-Pre | M-F1 | #failed | Acc | M-Rec | M-Pre | M-F1 | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.595 | 0.575 | 0.544 | 0.516 | 0 | 0.556 | 0.586 | 0.544 | 0.510 | 0 | ||
| Gemini Pro | 0.609 | 0.521 | 0.453 | 0.470 | 2 | 0.572 | 0.511 | 0.444 | 0.454 | 1 | |||
| Claude 2.1 | 0.375 | 0.510 | 0.474 | 0.415 | 2 | 0.328 | 0.466 | 0.447 | 0.369 | 1 | |||
| Llama-2 13B-chat | 0.406 | 0.188 | 0.191 | 0.188 | 0 | 0.384 | 0.179 | 0.180 | 0.178 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.633 | 0.532 | 0.551 | 0.470 | 0 | 0.594 | 0.531 | 0.494 | 0.438 | 0 | |||
| E-commerce LLM | EcomGPT | 0.191 | 0.362 | 0.341 | 0.188 | 6 | 0.196 | 0.375 | 0.336 | 0.178 | 13 | ||
| SoTA task-specific model | BERTweet | 0.733 | 0.503 | 0.530 | 0.511 | 0 | 0.729 | 0.507 | 0.524 | 0.513 | 0 | ||
| DeBERTaV3 | 0.768 | 0.567 | 0.607 | 0.573 | 0 | 0.764 | 0.565 | 0.591 | 0.567 | 0 | |||
| P5 | 0.611 | 0.199 | 0.157 | 0.156 | 0 | 0.620 | 0.200 | 0.124 | 0.153 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.783 | 0.619 | 0.618 | 0.612 | 0 | 0.770 | 0.604 | 0.601 | 0.600 | 0 | ||
| Llama-2 13B-chat | 0.791 | 0.616 | 0.641 | 0.616 | 0 | 0.781 | 0.627 | 0.645 | 0.629 | 0 | |||
| Llama-2 7B-chat | 0.790 | 0.620 | 0.652 | 0.634 | 0 | 0.769 | 0.583 | 0.599 | 0.589 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.801 | 0.643 | 0.676 | 0.655 | 0 | 0.789 | 0.619 | 0.650 | 0.632 | 0 | |||
| Flan-T5 XL | 0.771 | 0.645 | 0.638 | 0.620 | 0 | 0.743 | 0.594 | 0.592 | 0.582 | 0 | |||
| Phi-2 | 0.779 | 0.611 | 0.618 | 0.608 | 0 | 0.754 | 0.576 | 0.594 | 0.583 | 0 | |||
| Generalist | Flan-T5 XXL | 0.797 | 0.629 | 0.646 | 0.628 | 0 | 0.787 | 0.619 | 0.624 | 0.619 | 0 | ||
| Llama-2 13B-chat | 0.796 | 0.641 | 0.661 | 0.648 | 0 | 0.785 | 0.621 | 0.638 | 0.629 | 0 | |||
| Llama-2 7B-chat | 0.768 | 0.579 | 0.589 | 0.580 | 0 | 0.776 | 0.599 | 0.626 | 0.606 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.781 | 0.630 | 0.654 | 0.639 | 0 | 0.784 | 0.630 | 0.653 | 0.640 | 0 | |||
| Flan-T5 XL | 0.782 | 0.654 | 0.655 | 0.648 | 0 | 0.753 | 0.604 | 0.598 | 0.598 | 0 | |||
| Phi-2 | 0.780 | 0.588 | 0.619 | 0.596 | 0 | 0.758 | 0.552 | 0.590 | 0.565 | 0 | |||
@@
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric βAccβ, βM-Recβ, βM-Preβ, and βM-F1β represents accuracy, macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
@@
In this table, βINDβ indicates in-domain evaluation; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metric βM-Recβ, βM-Preβ, and βM-F1β represents macro recall, macro precision, and macro F1, respectively. On each task, the best performance is in bold.
@@
In this table, βINDβ indicates in-domain evaluation; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; βNPRβ is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||||||
| Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | Accuracy | Recall | Precision | F1 | Specificity | NPR | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.623 | 0.550 | 0.791 | 0.649 | 0.749 | 0.491 | 5 | 0.641 | 0.555 | 0.876 | 0.680 | 0.828 | 0.460 | 2 | ||
| Gemini Pro | 0.542 | 0.371 | 0.797 | 0.506 | 0.837 | 0.435 | 13 | 0.543 | 0.410 | 0.844 | 0.552 | 0.834 | 0.393 | 6 | |||
| Claude 2.1 | 0.424 | 0.177 | 0.671 | 0.280 | 0.850 | 0.375 | 90 | 0.384 | 0.146 | 0.769 | 0.245 | 0.904 | 0.326 | 73 | |||
| Llama-2 13B-chat | 0.534 | 0.608 | 0.638 | 0.623 | 0.406 | 0.375 | 0 | 0.541 | 0.605 | 0.688 | 0.644 | 0.401 | 0.317 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.522 | 0.539 | 0.647 | 0.588 | 0.493 | 0.383 | 120 | 0.537 | 0.523 | 0.725 | 0.608 | 0.567 | 0.352 | 126 | |||
| E-commerce LLM | EcomGPT | 0.318 | 0.051 | 0.283 | 0.086 | 0.779 | 0.322 | 254 | 0.286 | 0.085 | 0.403 | 0.140 | 0.726 | 0.266 | 245 | ||
| SoTA task-specific model | BERT | 0.749 | 0.970 | 0.726 | 0.830 | 0.368 | 0.877 | 0 | 0.803 | 0.831 | 0.876 | 0.853 | 0.742 | 0.668 | 0 | ||
| DeBERTaV3 | 0.504 | 0.310 | 0.769 | 0.441 | 0.839 | 0.413 | 0 | 0.501 | 0.345 | 0.826 | 0.487 | 0.841 | 0.370 | 0 | |||
| Task-specific | Flan-T5 XXL | 0.749 | 0.875 | 0.763 | 0.815 | 0.531 | 0.712 | 0 | 0.799 | 0.889 | 0.830 | 0.859 | 0.602 | 0.713 | 0 | ||
| Llama-2 13B-chat | 0.801 | 0.919 | 0.797 | 0.854 | 0.597 | 0.811 | 0 | 0.838 | 0.924 | 0.852 | 0.887 | 0.650 | 0.797 | 0 | |||
| Llama-2 7B-chat | 0.741 | 0.889 | 0.749 | 0.813 | 0.485 | 0.718 | 0 | 0.761 | 0.864 | 0.802 | 0.832 | 0.535 | 0.644 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.821 | 0.896 | 0.834 | 0.864 | 0.692 | 0.794 | 0 | 0.835 | 0.891 | 0.872 | 0.881 | 0.713 | 0.749 | 0 | |||
| Flan-T5 XL | 0.693 | 0.684 | 0.802 | 0.738 | 0.708 | 0.565 | 0 | 0.707 | 0.682 | 0.862 | 0.762 | 0.761 | 0.523 | 0 | |||
| Phi-2 | 0.765 | 0.942 | 0.751 | 0.835 | 0.460 | 0.820 | 0 | 0.781 | 0.950 | 0.779 | 0.856 | 0.411 | 0.791 | 0 | |||
| Generalist | Flan-T5 XXL | 0.774 | 0.910 | 0.773 | 0.836 | 0.540 | 0.776 | 0 | 0.814 | 0.914 | 0.832 | 0.871 | 0.596 | 0.760 | 0 | ||
| Llama-2 13B-chat | 0.808 | 0.864 | 0.838 | 0.851 | 0.711 | 0.752 | 0 | 0.818 | 0.862 | 0.872 | 0.867 | 0.723 | 0.705 | 0 | |||
| Llama-2 7B-chat | 0.817 | 0.921 | 0.814 | 0.864 | 0.638 | 0.824 | 0 | 0.836 | 0.931 | 0.845 | 0.886 | 0.627 | 0.807 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.797 | 0.880 | 0.814 | 0.846 | 0.654 | 0.759 | 0 | 0.832 | 0.885 | 0.872 | 0.878 | 0.717 | 0.740 | 0 | |||
| Flan-T5 XL | 0.765 | 0.888 | 0.774 | 0.827 | 0.553 | 0.741 | 0 | 0.819 | 0.891 | 0.852 | 0.871 | 0.662 | 0.735 | 0 | |||
| Phi-2 | 0.794 | 0.897 | 0.801 | 0.846 | 0.616 | 0.777 | 0 | 0.823 | 0.937 | 0.828 | 0.879 | 0.573 | 0.807 | 0 | |||
@@
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output; βNPRβ is the negative prediction rate. On each task, the best performance is in bold.
| Model | IND | OOD | |||||||||||
| R | P | F | BLEURT | #failed | R | P | F | BLEURT | #failed | ||||
| General-purpose LLMs | GPT-4 Turbo | 0.847 | 0.869 | 0.858 | 0.280 | 0 | 0.852 | 0.868 | 0.860 | 0.283 | 0 | ||
| Gemini Pro | 0.844 | 0.867 | 0.855 | 0.269 | 0 | 0.847 | 0.866 | 0.856 | 0.264 | 0 | |||
| Claude 2.1 | 0.848 | 0.835 | 0.841 | 0.314 | 0 | 0.851 | 0.833 | 0.842 | 0.325 | 0 | |||
| Llama-2 13B-chat | 0.845 | 0.780 | 0.811 | 0.261 | 0 | 0.845 | 0.775 | 0.808 | 0.260 | 0 | |||
| Mistral-7B-Instruct-v0.2 | 0.850 | 0.856 | 0.853 | 0.288 | 0 | 0.852 | 0.851 | 0.851 | 0.290 | 0 | |||
| E-commerce LLM | EcomGPT | 0.675 | 0.665 | 0.669 | 0.290 | 0 | 0.729 | 0.716 | 0.722 | 0.296 | 0 | ||
| SoTA task-specific model | GPT-4 Turbo | 0.847 | 0.869 | 0.858 | 0.280 | 0 | 0.852 | 0.868 | 0.860 | 0.283 | 0 | ||
| Task-specific | Flan-T5 XXL | 0.822 | 0.864 | 0.842 | 0.310 | 0 | 0.824 | 0.865 | 0.843 | 0.302 | 0 | ||
| Llama-2 13B-chat | 0.824 | 0.861 | 0.841 | 0.309 | 0 | 0.821 | 0.860 | 0.840 | 0.289 | 0 | |||
| Llama-2 7B-chat | 0.822 | 0.861 | 0.841 | 0.301 | 0 | 0.820 | 0.861 | 0.840 | 0.289 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.823 | 0.860 | 0.841 | 0.310 | 0 | 0.823 | 0.861 | 0.842 | 0.298 | 0 | |||
| Flan-T5 XL | 0.823 | 0.864 | 0.843 | 0.320 | 0 | 0.824 | 0.864 | 0.843 | 0.307 | 0 | |||
| Phi-2 | 0.817 | 0.855 | 0.835 | 0.283 | 0 | 0.817 | 0.856 | 0.835 | 0.270 | 0 | |||
| Generalist | Flan-T5 XXL | 0.824 | 0.865 | 0.844 | 0.224 | 0 | 0.823 | 0.864 | 0.843 | 0.206 | 0 | ||
| Llama-2 13B-chat | 0.823 | 0.861 | 0.841 | 0.215 | 0 | 0.822 | 0.861 | 0.841 | 0.195 | 0 | |||
| Llama-2 7B-chat | 0.822 | 0.860 | 0.840 | 0.208 | 0 | 0.819 | 0.859 | 0.838 | 0.188 | 0 | |||
| Mistral-7B Instruct-v0.2 | 0.822 | 0.864 | 0.842 | 0.213 | 0 | 0.821 | 0.862 | 0.840 | 0.194 | 0 | |||
| Flan-T5 XL | 0.823 | 0.864 | 0.843 | 0.227 | 0 | 0.824 | 0.865 | 0.844 | 0.211 | 0 | |||
| Phi-2 | 0.823 | 0.861 | 0.842 | 0.222 | 0 | 0.821 | 0.859 | 0.840 | 0.198 | 0 | |||
@@
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metrics βRβ, βPβ, βFβ and βBLEURTβ are detailed in TableΒ A1. On each task, the best performance is in bold. Note that we use GPT-4 Turbo as the SoTA task-specific model in this task.
| Model | ||||||||||
| IND | OOD | IND | ||||||||
| HR@1 | #failed | HR@1 | #failed | NDCG | #failed | |||||
| General-purpose LLMs | GPT-4 Turbo | 0.387 | 0 | 0.198 | 0 | 0.875 | 14 | |||
| Gemini Pro | 0.269 | 2 | 0.116 | 3 | 0.821 | 52 | ||||
| Claude 2.1 | 0.066 | 34 | 0.036 | 42 | 0.821 | 26 | ||||
| Llama-2 13B-chat | 0.056 | 0 | 0.050 | 0 | 0.815 | 0 | ||||
| Mistral-7B-Instruct-v0.2 | 0.164 | 1 | 0.108 | 0 | 0.842 | 4 | ||||
| E-commerce LLM | EcomGPT | 0.042 | 344 | 0.023 | 391 | 0.000 | 1000 | |||
| SoTA task-specific model | gSASRec / BERT | 0.249 | 0 | 0.065 | 0 | 0.839 | 0 | |||
| Recformer / DeBERTaV3 | 0.265 | 0 | 0.081 | 0 | 0.859 | 0 | ||||
| Task-specific | Flan-T5 XXL | 0.467 | 0 | 0.252 | 0 | 0.881 | 0 | |||
| Llama-2 13B-chat | 0.518 | 0 | 0.263 | 0 | 0.879 | 0 | ||||
| Llama-2 7B-chat | 0.517 | 0 | 0.228 | 0 | 0.867 | 0 | ||||
| Mistral-7B Instruct-v0.2 | 0.535 | 0 | 0.268 | 0 | 0.883 | 0 | ||||
| Flan-T5 XL | 0.436 | 0 | 0.226 | 0 | 0.875 | 0 | ||||
| Phi-2 | 0.413 | 5 | 0.219 | 10 | 0.858 | 0 | ||||
| Generalist | Flan-T5 XXL | 0.512 | 0 | 0.262 | 0 | 0.885 | 0 | |||
| Llama-2 13B-chat | 0.526 | 0 | 0.273 | 0 | 0.870 | 0 | ||||
| Llama-2 7B-chat | 0.517 | 0 | 0.261 | 0 | 0.868 | 0 | ||||
| Mistral-7B Instruct-v0.2 | 0.542 | 0 | 0.280 | 0 | 0.876 | 0 | ||||
| Flan-T5 XL | 0.463 | 0 | 0.256 | 0 | 0.868 | 0 | ||||
| Phi-2 | 0.479 | 5 | 0.241 | 8 | 0.870 | 0 | ||||
@@
In this table, βINDβ and βOODβ indicates in-domain evaluation and out-of-domain evaluation, respectively; βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together; #failed refers to the number of failure cases that we cannot extract meaningful results from the model output. The metrics βHR@1β and βNDCGβ are detailed in AppendixΒ A. On each task, the best performance is in bold.
Appendix H OOD Evaluation